OpenStudy (anonymous):

An acute angle θ is in a right triangle with sin θ = 6/7 . What is the value of cot θ

Parth (parthkohli):

What have you tried till now?

Parth (parthkohli):

Well,\[\tan \theta = {opposite \over adjacent}\]From \(\sin \theta\), we know that "opposite" is 6.

Parth (parthkohli):

Also, you must must find the adjacent side through Pythagorean Theorem.\[6^2 + b^2 = 7^2\]

Parth (parthkohli):

Yeah, also recall that\[\cot \theta = {adjacent \over opposite}\]

OpenStudy (anonymous):

so thats right? :)

Parth (parthkohli):

\[b = \sqrt{7^2 - 6^2} = \sqrt{13}\]Yay!

OpenStudy (anonymous):

ohh so 6/√13 ?

Parth (parthkohli):

No, your answer is right.

OpenStudy (anonymous):

ok i gt confused for a seccc

OpenStudy (anonymous):

love ur pic btw